OpenStudy (samsan9):

combine into a single logarithm 3log(x+y)+2log(x-y)-log(x^2+2y)=

OpenStudy (lasttccasey):

Well there are a few log rules that will help you solve this. Review this site for the 5 rules if you are unfamiliar at all with them: http://www.mathwords.com/l/logarithm_rules.htm

OpenStudy (anonymous):

log a - log b = log (a/b) and log a + log b = log (ab) and n log a = log (a^n)

OpenStudy (samsan9):

so the part after the -log is in fraction form?

OpenStudy (anonymous):

yup

OpenStudy (samsan9):

ok so now what?

OpenStudy (anonymous):

oh boy, I am lazy and tired. I want to sleep , hihihihi. why were they all offline. \(3 log (x+y) = log (x+y)^3\) \(2log(x-y)= log (x-y)^2\) now combine follow the rule above \( log (x+y)^3 +log(x-y)^2 -log(x^2+2y) = log \dfrac{(x+y)^3(x-y)^2}{ x^2+2y}\)